OpenStudy (candycanegone):

Factor. 2y(4a + b) − (b + 4a)

OpenStudy (candycanegone):

@alekos i got (2y+1)(b+4a)

OpenStudy (alekos):

Nice try. Its actually (2y-1)(b+4a)

OpenStudy (candycanegone):

why is it minus?

OpenStudy (alekos):

because the second term is - (b+4a)

OpenStudy (candycanegone):

oh i thought we changed it to plus when we change the variable around to match

OpenStudy (alekos):

No, you must carry through the minus

OpenStudy (candycanegone):

can i show you what i mean . this person had factor. c(a-2) -b(2-a) and got c(a-2) + b(a-2)

OpenStudy (candycanegone):

the got (a-2) and c+b)

OpenStudy (candycanegone):

it it because there was a b there?

OpenStudy (alekos):

no its because -b(2-a) was changed to +b(a-2)

OpenStudy (candycanegone):

can i ask why it was changed like what is the rule?

OpenStudy (alekos):

oh ok. -b(2-a) = (-1)b(2-a) = b(-1)(2-a) = b[(2)(-1)-a(-1)] = b(-2+a) = b(a-2)

OpenStudy (alekos):

can you follow that?

OpenStudy (candycanegone):

ummm lol

OpenStudy (alekos):

that ok?

OpenStudy (candycanegone):

can you put it in words like there is a - infront of parenthesis we keep the negative sign

OpenStudy (alekos):

I'll let someone else give it to you in words

OpenStudy (karim728):